# Effects of boundary conditions on quantum nanoresonators: decoherence-free subspaces

### v2

@@ I Introduction
Due to the large number of particles, modeling a nanoresonator or a nanobeam is not a simple task. Classical models limit their treatment as a continuous medium, and discrete aspects of the material’s lattice may be relevant. Among the approaches is atomic modeling, which essentially involves solving the Schrödinger equation for a system of $N$ atoms. Atomic modeling is usually divided into molecular dynamics, Monte Carlo, density functional theory, tight bonding molecular dynamics, local density, Morse potential function model, and ab initio approaches [44].
@@ II Classical beam modeling
We briefly summarize the main results from the Euler-Bernoulli beam theory, with detailed discussions available in standard textbooks, such as [5]. We assume that the beam’s displacements are confined to the $xz$-plane and denote them by $w(x,t)$, where positive $w$ indicates downward displacement. The potential energy of the beam is given by
@@
where $E$ is the Young’s modulus of the material, and $I$ represents the area moment of inertia of the beam’s cross-section about the $y$-axis. The kinetic energy of the beam is
@@
where $\rho$ is the mass density of the beam and $A$ is the cross-sectional area of the beam about the $y$-axis.
@@ III.2 Effects of vacuum: The Phonon Casimir Effect
Note from (32) that $F\propto L^{-3}$, and it is proportional to the energy fundamental state of the first mode, $k=1$, divided by the volume of the beam. A discussion of a similar phonon Casimir effect can be found in references [28, 32]. Consequently, it is very small and attractive. In the next calculations, we will neglect this force. On the other hand, if the term $AL$ is small and $F$ isn’t negligible, the force $F$ should be considered from the beginning in the Euler-Bernoulli modeling [21].
@@ IV.0.1 Some considerations about a more realistic thermal bath
If we consider a zero-temperature thermal bath [9], we have a decay rate that depends on $D(\omega)d\omega$, which represents the number of modes in the frequency interval between $\omega$ and $\omega+d\omega$ of the thermal bath. Since the modes represent vibrational states of the same beam, the density of states of the thermal bath is the same. However, since different vibration modes represent deformations with different wavelengths, it is reasonable to imagine that the coupling between the baths must be different for each vibrational mode, and therefore the rates of loss of coherence ($\kappa_{i}$) can be very different, where $\kappa_{i}$ is the decay rate of the vibrational mode $i$. If we assume that the decay rate is $\kappa=\eta\omega$ as shown in ref. [55], then we can have $\kappa_{A}\ll\kappa_{B}$. If the modes were uncoupled, the mode $B$ would lose its coherence much faster than mode $A$. Now, including a coupling, it makes them decay at a rate $\kappa_{AB}$, and we have $\kappa_{AB}\approx\kappa_{A}$, which is a necessary condition for the existence of a decoherence-free subspace [9]. In this case, it is not a subspace without decoherence but one with a significantly longer half-life.

### v3

@@ I Introduction
Due to the large number of particles, modeling a nanoresonator or a nanobeam is not a simple task. Classical models limit their treatment as a continuous medium, and discrete aspects of the material’s lattice may be relevant. Among the approaches is atomic modeling, which essentially involves solving the Schrödinger equation for a system of NN atoms. Atomic modeling is usually divided into molecular dynamics, Monte Carlo, density functional theory, tight bonding molecular dynamics, local density, Morse potential function model, and ab initio approaches [44].
@@ II Classical beam modeling
We briefly summarize the main results from the Euler-Bernoulli beam theory, with detailed discussions available in standard textbooks, such as [5]. We assume that the beam’s displacements are confined to the xzxz-plane and denote them by $w(x,t)$, where positive $w$ indicates downward displacement. The potential energy of the beam is given by
@@
where $E$ is the Young’s modulus of the material, and II represents the area moment of inertia of the beam’s cross-section about the $y$-axis. The kinetic energy of the beam is
@@
where $\rho$ is the mass density of the beam and AA is the cross-sectional area of the beam about the $y$-axis.
@@ III.2 Effects of vacuum: The Phonon Casimir Effect
Note from (32) that $F\propto L^{-3}$, and it is proportional to the energy fundamental state of the first mode, $k=1$, divided by the volume of the beam. A discussion of a similar phonon Casimir effect can be found in references [28, 32]. Consequently, it is very small and attractive. In the next calculations, we will neglect this force. On the other hand, if the term ALAL is small and $F$ isn’t negligible, the force $F$ should be considered from the beginning in the Euler-Bernoulli modeling [21].
@@ IV.0.1 Some considerations about a more realistic thermal bath
If we consider a zero-temperature thermal bath [9], we have a decay rate that depends on $D(\omega)d\omega$, which represents the number of modes in the frequency interval between $\omega$ and $\omega+d\omega$ of the thermal bath. Since the modes represent vibrational states of the same beam, the density of states of the thermal bath is the same. However, since different vibration modes represent deformations with different wavelengths, it is reasonable to imagine that the coupling between the baths must be different for each vibrational mode, and therefore the rates of loss of coherence ($\kappa_{i}$) can be very different, where $\kappa_{i}$ is the decay rate of the vibrational mode ii. If we assume that the decay rate is $\kappa=\eta\omega$ as shown in ref. [55], then we can have $\kappa_{A}\ll\kappa_{B}$. If the modes were uncoupled, the mode $B$ would lose its coherence much faster than mode AA. Now, including a coupling, it makes them decay at a rate $\kappa_{AB}$, and we have $\kappa_{AB}\approx\kappa_{A}$, which is a necessary condition for the existence of a decoherence-free subspace [9]. In this case, it is not a subspace without decoherence but one with a significantly longer half-life.